# Dissipative currents and transport coefficients in relativistic spin hydrodynamics

### v1

@@ VI Summary and outlook
With this foundation, one can formulate the evolution equations for the energy-momentum tensor, conserved vector current and spin tensor at first order in gradients, commonly referred to as the relativistic Navier-Stokes limit. However, the presence of 23 transport coefficients, which have to be either extracted from a suitable microscopic theory or fitted to data, may pose considerable challenges for future numerical studies, unless some of the gradients are neglected.
@@ Appendix A Decomposition of tensors into irreducible components under the rotation group
A four-vector $V^{\mu}$ can be decomposed into two irreducible components under the rotation group associated to the space orthogonal to $u$: a scalar, obtained by contracting $V$ with $u$ and a vector with vanishing component along $u$. We will denote this decomposition of $V$ onto the irreducible representations $j=0$ and $j=1$ of $\mathrm{SO}(3,\mathbb{R})$ that is scalar and the vector representations - by writing:
@@
where the extra indices $\alpha$ and $\beta$ again signify that tose expressions are projectors (or proportional to projectors) to be applied to tensors. The $2$ in eq. (68) denotes the irreducible 5-dimensional representation pertaining to Euclidean symmetric traceless tensors. Note that the anti-symmetric tensor requires the use of the Levi-Civita tensor to generate an independent vector projector.
By using the above decompositions as building blocks, we can decompose tensors of any rank into irreducible components under the rotation group. For example, let us consider the gradient of the spin potential $\partial_{\rho}\Omega_{\lambda\nu}$, which is a rank-3 tensor antisymmetric in the last 2 indices. We can then write, by using the decompositions in (67) and (68):
@@
where the first factor $(0\oplus 1)$ refers to the gradient vector $\partial_{\rho}$ and the second, i.e. $(1\oplus 1)$, to the spin potential tensor $\Omega$. By expanding the above tensor product and taking into account the known rules for the decomposition into irreducible representations, we get:
@@
[…]

### v2

@@ VI Summary and outlook
With this foundation, one can formulate the evolution equations for the energy-momentum tensor, conserved vector current and spin tensor at first order in gradients, commonly referred to as the relativistic Navier-Stokes limit. However, the presence of 23 transport coefficients, which have to be either extracted from a suitable microscopic theory or fitted to data, may pose considerable challenges for future numerical studies, unless some of the gradients are neglected. Imposing the consequences of Onsager’s reciprocity theorem on our formalism may also relate some coefficients of different gradients.
@@ Appendix A Decomposition of tensors into irreducible components under the rotation group
A four-vector $V^{\mu}$ can be decomposed into two irreducible components under the rotation group associated to the space orthogonal to $u$: a scalar, obtained by contracting $V$ with $u$ and a vector with vanishing component along $u$. We will denote this decomposition of $V$ onto the irreducible representations $j=0$ and $j=1$ of $\mathrm{SO}(3,\mathbb{R})$ - that is scalar and the vector representations - by writing:
@@
where the extra indices $\alpha$ and $\beta$ again signify that those expressions are projectors (or proportional to projectors) to be applied to tensors. The $2$ in eq. (68) denotes the irreducible 5-dimensional representation pertaining to Euclidean symmetric traceless tensors. Note that the anti-symmetric tensor requires the use of the Levi-Civita tensor to generate an independent vector projector.
By using the above decompositions as building blocks, we can decompose tensors of any rank into irreducible components under the rotation group. For example, let us consider the gradient of the reduced spin potential $\partial_{\rho}\Omega_{\lambda\nu}$, which is a rank-3 tensor antisymmetric in the last 2 indices. We can then write, by using the decompositions in (67) and (68):
@@
where the first factor $(0\oplus 1)$ refers to the gradient vector $\partial_{\rho}$ and the second, i.e., $(1\oplus 1)$, to the reduced spin potential tensor $\Omega$. By expanding the above tensor product and taking into account the known rules for the decomposition into irreducible representations, we get:
@@
$0\,\otimes\,0=0\qquad 0\,\otimes\,1=1\qquad 1\,\otimes\,1=(0,1,2)\qquad 1\,\otimes\,2=(1,2,3)\qquad 2\,\otimes\,2=(0,1,2,3,4)$ (73)